# The Honeycomb Conjecture in normed planes and an alpha-convex variant of a theorem of Dowker

### v1

@@ Abstract.
The Honeycomb Conjecture states that among tilings with unit area cells in the Euclidean plane, the average perimeter of a cell is minimal for a regular hexagonal tiling. This conjecture was proved by L. Fejes Tóth for convex tilings, and by Hales for not necessarily convex tilings. In this paper we investigate the same question for tilings of a given normed plane, and show that among normal, convex tilings in a normed plane, the average squared perimeter of a cell is minimal for a tiling whose cells are translates of a centrally symmetric hexagon. We also show that the question whether the same statement is true for the average perimeter of a cell is closely related to an $\alpha$-convex variant of a theorem of Dowker on the area of polygons circumscribed about a convex disk. Exploring this connection we find families of norms in which the average perimeter of a cell of a tiling is minimal for a hexagonal tiling, and prove some additonal related results. Finally, we apply our method to give a partial answer to a problem of Steinhaus about the isoperimetric ratios of cells of certain tilings in the Euclidean plane, appeared in an open problem book of Croft, Falconer and Guy.
@@
Partially supported by the National Research, Development and Innovation Office, NKFI, K-147544 grant, and BME Water Sciences & Disaster Prevention TKP2020 Institution Excellence Subprogram, grant no. TKP2020 BME-IKA-VIZ
@@ 1. Introduction
One of the classic problems of geometry, regarding planar mosaics, is the Honeycomb Conjecture, which states, roughly speaking, that in a decomposition of the Euclidean plane into cells of equal area, the average perimeter of the cells is minimal for the regular hexagonal grid; i.e. a tiling in which every cell is a unit area regular hexagon. This conjecture first appeared in Roman time in a book of Varro about agriculture [26]. Despite its ancient origin, it took a surprisingly long time to find a satisfactory proof for this conjecture [18]: it was solved by L. Fejes Tóth for normal, convex mosaics [10] in the 1940s , while the most general version is due to Hales [15] in 2001, who dropped the condition of convexity. Despite its solution, this problem still seems to be interesting for mathematicians [3, 4], engineers [28] and even for philosophers [19].
Clearly, it is a natural question to ask if a similar statement holds for tilings in any normed plane. The aim of this paper to investigate this problem. To do it, we recall a few well-known facts about normed planes. In particular, we recall that every origin-symmetric convex disk $M$ (i.e. every compact, convex set $M$ with nonempty interior and satisfying $M=-M$) is the unit disk of a normed plane, and the unit disk of a normed plane is an origin-symmetric convex disk. In the paper, for a normed plane $\mathcal{M}$, we denote the unit disk of $\mathcal{M}$ by $M$, and the norm of a point $p\in\mathcal{M}$ by $||p||_{M}$. For convenience, we imagine that $\mathcal{M}$ also have an underlying Euclidean structure, and define the Euclidean norm of $p$ by $||p||$.
@@
Is it true that in any normal, convex tiling in any normed pane $\mathcal{M}$, with unit area cells, the average perimeter of a cell, if it is exists, is minimal for a hexagonal tiling?
@@
Let $\mathcal{T}$ be a convex, normal tiling in the normed plane $\mathcal{M}$. For any $R>0$, let $\mathcal{T}(R)=$ denote the family of cells of $\mathcal{T}$ contained in $R\mathbf{B}^{2}$. Let $\alpha>0$. We define the lower average $\alpha$th powered perimeter of a cell of $\mathcal{T}$ as the quantity
@@
For any $\alpha,\beta\in[0,\infty)$ with $\alpha<\beta$ and any normal, convex tiling in $\mathcal{M}$, we have $\underline{P}_{\alpha}(\mathcal{T})\leq\underline{P}_{\beta}(\mathcal{T})$ and $\overline{P}_{\alpha}(\mathcal{T})\leq\overline{P}_{\beta}(\mathcal{T})$. Furthermore, if $\mathcal{T}$ is a hexagonal tiling, we have equality in all the previous inequalities.
Our first result is the following.
@@
(1.2) $\frac{n-6}{n-m}A_{K}^{\alpha}(m)+\frac{6-m}{n-m}A_{K}^{\alpha}(n)\geq 2A_{K}^{\alpha}(6)$
@@
It is an elementary exercise to check $A_{\mathbf{B}^{2}}(n)=n\tan\frac{\pi}{n}$, implying that $\mathbf{B}^{2}$ satisfies the log-Dowker property. Thus, Theorem 1.9 immediately yields Corollary 1.10.
@@
In Section 2, we prove Theorems 1.4 and 1.9. In Sections 3 and 4 we investigate the $\alpha$-Dowker and $\alpha$-convex properties of convex disks and normed planes, respectively. More specifically, in Section 3 we consider convex polygons and polygonal norms, while in Section 4 we examine properties of convex disks and norms which are not necessarily polygonal. In Section 5 we consider a problem of Steinhaus, appeared in the problem book [7] of Croft, Falconer and Guy, asking if it is true that among tilings of the Euclidean plane with tiles whose diameters are at least some universal constant $D>0$, the maximum isoperimetric ratio of the cells is minimal for a regular hexagonal tiling. Using the approach of the proof of Theorem 1.4, we show that for convex, normal tilings, a stronger statement holds in any normed plane, and also investigate possible generalizations of our method for this problem. We finish the paper with an additional remark in Section 6.
In our investigation, we denote by $\int(X)$, $\operatorname{bd}(X)$ and $\operatorname{conv}(X)$ the interior, boundary and the convex hull of a set $X$, and for any two points $p,q$, we denote by $[p,q]$ the closed straight line segment connecting them. Finally, we denote the Euclidean unit circle $\operatorname{bd}\mathbf{B}^{2}$ by $\mathbb{S}^{1}$.
@@
First, we observe that Theorem 1.9, combined with Theorem 1.5 readily implies Theorem 1.4. Thus, we need only prove Theorem 1.9. Throughout this section, we deal with a given convex, normal tiling $\mathcal{T}$ in a normed plane $\mathcal{M}$. We permit the angles of cells to be straight angles; due to this, we can regard $\mathcal{T}$ as an edge-to-edge tiling, in which every edge of every cell belongs to exactly one more cell. During our investigation, $\mathcal{T}(R)$ denotes the family of cells of $\mathcal{T}$ contained in $R\mathbf{B}^{2}$, and for every $C\in\mathcal{T}$, $v(C)$ denotes the number of sides of $C$. We start with collecting some elementary observation and known facts about tilings and normed planes.
@@
For every cell $C\in\mathcal{T}$, there is a Euclidean disk of radius $3R$ that contains $C$ as well as all its neighbors. The area of every cell is at least $r^{2}\pi$. Thus, the assertion follows from a simple area estimate. ∎
@@
Since the diameter of every cell $C\in\mathcal{T}$ is at most $2\hat{R}$, every cell is contained in the closure of the set $(R+2\hat{R})\mathbf{B}^{2}\setminus(R-2\hat{R})\mathbf{B}^{2}$. Thus, $\operatorname{card}(\mathcal{T}_{bd}(R))\leq\frac{(R+2\hat{R})^{2}\pi-(R-2\hat{R})^{2}\pi}{\hat{r}^{2}\pi}=8\hat{R}R$, implying that $\operatorname{card}(\mathcal{T}_{bd}(R))=\mathcal{O}(R)$. On the other hand, the union of these cells covers $R\mathbb{S}^{2}$, and the diameter of each cell is at most $2\hat{R}$. Thus, $\operatorname{card}(\mathcal{T}_{bd}(R))\geq\frac{2\pi}{2\arcsin\frac{\hat{R}}{R}}$, implying that $\operatorname{card}(\mathcal{T}_{bd}(R))=\Omega(R)$. ∎
@@
For any $R>2\bar{R}$, set $v_{R}(\mathcal{T})=\frac{\sum_{C\in\mathcal{T}(R)}v(C)}{\operatorname{card}(\mathcal{T}(R))}$. Then, $v_{R}(\mathcal{T})\leq 6+\mathcal{O}\left(\frac{1}{R}\right)$, and $\overline{v}(\mathcal{T})\leq 6$.
@@
Consider the graph $G$ whose vertices are the vertices of the cells of $\mathcal{T}(R)$, and two vertices are connected by an edge if they are the endpoints of an edge of a cell in $\mathcal{T}(R)$. Then $G$ is a planar graph. Let $V(G)$, $E(G)$ and $F(G)$ denote the numbers of vertices, edges and faces of $G$. Furthermore, let $\mathcal{E}_{bd}(G)$ denote the family of the edges of $G$ in the boundary of $\bigcup\mathcal{T}(R)$. We note that every edge in $\mathcal{E}_{bd}(G)$ belongs to a cell of $\mathcal{T}$ intersecting $R\mathbb{S}^{1}$, and thus, it follows from Lemmas 2.1 and 2.2 that $E_{bd}(G):=\operatorname{card}\operatorname{card}(\mathcal{E}_{bd}(G))=\Theta(R)$. These lemmas, an area estimate and the observation that $\bigcup\mathcal{T}(R)$ contains $(R-2\bar{R})\mathbf{B}^{2}$ and is contained in $(R+2\bar{R})\mathbf{B}^{2}$ shows that $V(G)$, $E(G)$ and $F(G)$ are all of order of magnitude $\Theta(R^{2})$.
We observe that there might be vertices of $G$ whose degree is less than $3$. If $p$ is such a vertex, then its degree is $2$, it lies in $\operatorname{bd}\left(\bigcup\mathcal{T}(R)\right)$, and it belongs to two edges in $\mathcal{E}_{bd}(G)$. Thus, the number of such edges is at most $E_{bd}(G)=\mathcal{O}(R)$. We apply series reduction on these edges, and obtain another planar graph $G^{\prime}$ in which the degree of every vertex is at least $3$. If $V(G^{\prime}),E(G^{\prime})$ and $F(G^{\prime})$ denotes that numbers of vertices, edges and faces of $G^{\prime}$, then by the construction of $G^{\prime}$ we have $V(G^{\prime})\leq V(G)$ and $E(G^{\prime})\leq E(G)$ with $V(G)-V(G^{\prime})=E(G)-E(G^{\prime})=\mathcal{O}(R)$, and $F(G^{\prime})=F(G)$.
Applying Euler’s formula for $G^{\prime}$ and using the fact that the degree of every vertex of $G^{\prime}$ is $3$ yields that $E(G^{\prime})\leq 3F(G^{\prime})-6$, implying that $E(G)\leq 3F(G)+\mathcal{O}(R)$. On the other hand, $v_{R}(\mathcal{T})\leq\frac{2E(G)}{F(G)}$, and as $E(G)$ and $F(G)$ are of $\Theta(R^{2})$, it follows that $v_{R}(\mathcal{T})\leq 6+\mathcal{O}\left(\frac{1}{R}\right)$, which readily yields the assertion. ∎
@@
Figure 1. The isoperimetrix $M_{iso}$ of a norm with unit disk $M$. The dotted circle is the Euclidean unit disk $\mathbf{B}^{2}$ centered at $o$. The left-hand side panel shows $M$ and its polar $M^{\circ}$, the isoperimetrix in the righ-hand side panel is a rotated copy of $M^{\circ}$ rotated by $\frac{\pi}{2}$.
@@
A variant of this inequality, in Euclidean plane, states that among convex polygons with given perimeter and fixed directions of its sides, the ones circumscribed about a circular disk have maximal area. This statement is called L’huilier’s inequality, which is generalized for arbitrary normed planes by Chakerian [6] in the following way.
@@
Now we are ready to prove Theorem 1.9. We simplicity, we assume that $M_{iso}$ satisfies the $\alpha$-Dowker property for some $\alpha\in(0,\infty)$; for the case $\alpha=0$ a straightforward modification of our argument can be applied. From now on we also assume that every cell of $\mathcal{T}$ has unit area. We need to consider the following quantity.
$\underline{P}_{2\alpha}(\mathcal{T})=\liminf_{R\to\infty}\frac{\sum_{C\in\mathcal{T}(R)}\left(\operatorname{perim}_{M}(C)\right)^{\alpha}}{\operatorname{card}(\mathcal{T}(R))}.$
@@
$\underline{P}_{2\alpha}(\mathcal{T})\geq 4\liminf_{R\to\infty}\frac{\sum_{C\in\mathcal{T}(R)}\left(\operatorname{area}(C^{*})\right)^{\alpha}}{\operatorname{card}\mathcal{T}(R)}\geq 4\lim_{R\to\infty}\frac{\sum_{C\in\mathcal{T}(R)}\left(A_{M_{iso}}(v(C))\right)^{\alpha}}{\operatorname{card}\mathcal{T}(R)}.$
where $A_{M_{iso}}(v(C))$ denotes the area of a minimum area polygon with $v(C)$ sides, circumscribed about $M_{iso}$ (see Theorem 1.5). By our assumptions, the sequence $\{\left(A_{M_{iso}}(n)\right)^{\alpha}\}$ is convex, and by Lemma 2.4 $\overline{v}(\mathcal{T})\leq 6$. Thus,
$\underline{P}_{2\alpha}(\mathcal{T})\geq 4\left(A_{M_{iso}}(6)\right)^{\alpha}.$
Now the only thing we need to show that for some hexagonal tling $\mathcal{T}_{hex}$ of $\mathcal{M}$, we have $P_{2\alpha}(\mathcal{T}_{hex})=4\left(A_{M_{iso}}(6)\right)^{\alpha}$. To do it we recall another theorem of Dowker from [8].
@@
Let $K$ by an $o$-symmetric plane convex body. Then, for every $m\geq 2$, there is a centrally symmetric convex $(2m)$-gon $P$ circumscribed about $K$ with $\operatorname{area}(P)=A_{K}(2m)$.
By Theorem 2.7, there is an $o$-symmetric hexagon $H$ circumscribed about $M_{iso}$ of area $A_{M_{iso}}(6)$. Let $\mathcal{T}_{hex}$ be a tiling whose cells are translates of $H$. Then the desired equality follows from the equality part of Theorem 2.6.

### v2

@@ Abstract.
The Honeycomb Conjecture states that among tilings with unit area cells in the Euclidean plane, the average perimeter of a cell is minimal for a regular hexagonal tiling. This conjecture was proved by L. Fejes Tóth for convex tilings, and by Hales for not necessarily convex tilings. In this paper we investigate the same question for tilings of a given normed plane, and show that among normal, convex tilings in a normed plane, the average squared perimeter of a cell is minimal for a tiling whose cells are translates of a centrally symmetric hexagon. We also show that the question whether the same statement is true for the average perimeter of a cell is closely related to an $\alpha$-convex variant of a theorem of Dowker on the area of polygons circumscribed about a convex disk. Exploring this connection we find families of norms in which the average perimeter of a cell of a tiling is minimal for a hexagonal tiling, and prove some additional related results. Finally, we apply our method to give a partial answer to a problem of Steinhaus about the isoperimetric ratios of cells of certain tilings in the Euclidean plane, appeared in an open problem book of Croft, Falconer and Guy.
@@
Partially supported by the ERC Advanced Grant “ERMiD”, the National Research, Development and Innovation Office, NKFI, K-147544 grant, and BME Water Sciences & Disaster Prevention TKP2020 Institution Excellence Subprogram, grant no. TKP2020 BME-IKA-VIZ
@@ 1. Introduction
One of the classic problems of geometry, regarding planar mosaics, is the Honeycomb Conjecture, which states, roughly speaking, that in a decomposition of the Euclidean plane into cells of equal area, the average perimeter of the cells is minimal for the regular hexagonal grid; i.e. a tiling in which every cell is a unit area regular hexagon (for an investigation of mathematically rigorous variants of this problem, see [18]). This conjecture first appeared in Roman time in a book of Varro about agriculture [26]. Despite its ancient origin, it took a surprisingly long time to find a satisfactory proof for this conjecture: it was solved by L. Fejes Tóth for normal, convex mosaics [10] in the 1940s , while the most general version is due to Hales [15] in 2001, who dropped the condition of convexity; the proofs of both statements require heavy computation based on the metric properties of the Euclidean plane. Despite its solution, this problem still seems to be interesting for mathematicians [3, 4], engineers [28] and even for philosophers [19].
Clearly, it is a natural question to ask if a similar statement holds for tilings in any normed plane. The aim of this paper is to investigate this problem. To do it, we recall a few well-known facts about normed planes. In particular, we recall that every origin-symmetric convex disk $M$ (i.e. every compact, convex set $M$ with nonempty interior and satisfying $M=-M$) is the unit disk of a normed plane, and the unit disk of a normed plane is an origin-symmetric convex disk. In the paper, for a normed plane $\mathcal{M}$, we denote the unit disk of $\mathcal{M}$ by $M$, and the norm of a point $p\in\mathcal{M}$ by $||p||_{M}$. For convenience, we imagine that $\mathcal{M}$ also has an underlying Euclidean structure, and denote the Euclidean norm of $p$ by $||p||$.
@@
Is it true that in any normal, convex tiling in any normed pane $\mathcal{M}$, with unit area cells, the average perimeter of a cell, if it exists, is minimal for a hexagonal tiling?
@@
Let $\mathcal{T}$ be a convex, normal tiling in the normed plane $\mathcal{M}$. For any $R>0$, let $\mathcal{T}(R)$ denote the family of cells of $\mathcal{T}$ contained in $R\mathbf{B}^{2}$. Let $\alpha>0$. We define the lower average $\alpha$th powered perimeter of a cell of $\mathcal{T}$ as the quantity
@@
For any $\alpha,\beta\in(0,\infty)$ with $\alpha<\beta$ and any normal, convex tiling $\mathcal{T}$ in $\mathcal{M}$, we have $\exp(\underline{P}_{\log}(\mathcal{T}))\leq\left(\underline{P}_{\alpha}(\mathcal{T})\right)^{1/\alpha}\leq\left(\underline{P}_{\beta}(\mathcal{T})\right)^{1/\beta}$ and $\exp(\overline{P}_{\log}(\mathcal{T}))\leq\left(\overline{P}_{\alpha}(\mathcal{T})\right)^{1/\alpha}\leq\left(\overline{P}_{\beta}(\mathcal{T})\right)^{1/\beta}$. Furthermore, if $\mathcal{T}$ is a hexagonal tiling, we have equality in all the previous inequalities.
Our first result is the following theorem, which, as it is shown by Remark 1.3, can be regarded as a ‘weaker’ variant of the Honeycomb Conjecture.
@@
(1.2) $\frac{n-6}{n-m}A_{K}^{\alpha}(m)+\frac{6-m}{n-m}A_{K}^{\alpha}(n)\geq A_{K}^{\alpha}(6)$
@@
It is an elementary exercise to check that $A_{\mathbf{B}^{2}}(n)=n\tan\frac{\pi}{n}$, implying that $\mathbf{B}^{2}$ satisfies the log-Dowker property. Thus, Theorem 1.9 immediately yields Corollary 1.10.
@@
In Section 2, we prove Theorems 1.4 and 1.9. In Sections 3 and 4 we investigate the $\alpha$-Dowker and $\alpha$-honeycomb properties of convex disks and normed planes, respectively. More specifically, in Section 3 we consider convex polygons and polygonal norms, while in Section 4 we examine properties of convex disks and norms which are not necessarily polygonal. In Section 5 we consider a problem of Steinhaus, appeared in the problem book [7] of Croft, Falconer and Guy, asking if it is true that among tilings of the Euclidean plane with tiles whose diameters are at least some universal constant $D>0$, the maximum isoperimetric ratio of the cells is minimal for a regular hexagonal tiling. Using the approach of the proof of Theorem 1.4, we show that for convex, normal tilings, a stronger statement holds in any normed plane, and also investigate possible generalizations of our method for this problem. We finish the paper with an additional remark in Section 6.
In our investigation, we denote by $\operatorname{int}(X)$, $\operatorname{bd}(X)$ and $\operatorname{conv}(X)$ the interior, boundary and the convex hull of a set $X$, and for any two points $p,q$, we denote by $[p,q]$ the closed straight line segment connecting them. Finally, we denote the Euclidean unit circle $\operatorname{bd}\mathbf{B}^{2}$ by $\mathbb{S}^{1}$.
@@
First, we observe that Theorem 1.9, combined with Theorem 1.5 readily implies Theorem 1.4. Thus, we need only prove Theorem 1.9. Throughout this section, we deal with a given convex, normal tiling $\mathcal{T}$ in a normed plane $\mathcal{M}$. We permit the angles of cells to be straight angles; due to this, we can regard $\mathcal{T}$ as an edge-to-edge tiling, in which every edge of every cell belongs to exactly one more cell. During our investigation, $\mathcal{T}(R)$ denotes the family of cells of $\mathcal{T}$ contained in $R\mathbf{B}^{2}$, and for every $C\in\mathcal{T}$, $v(C)$ denotes the number of sides of $C$. We start with collecting some elementary observations and known facts about tilings and normed planes.
@@
For every cell $C\in\mathcal{T}$, there is a Euclidean disk of radius $3\hat{R}$ that contains $C$ as well as all its neighbors. The area of every cell is at least $\hat{r}^{2}\pi$. Thus, the assertion follows from a simple area estimate. ∎
@@
Since the diameter of every cell $C\in\mathcal{T}$ is at most $2\hat{R}$, every cell is contained in the closure of the set $(R+2\hat{R})\mathbf{B}^{2}\setminus(R-2\hat{R})\mathbf{B}^{2}$. Thus, $\operatorname{card}(\mathcal{T}_{bd}(R))\leq\frac{(R+2\hat{R})^{2}\pi-(R-2\hat{R})^{2}\pi}{\hat{r}^{2}\pi}=\frac{8\hat{R}R}{\hat{r}^{2}}$, implying that $\operatorname{card}(\mathcal{T}_{bd}(R))=\mathcal{O}(R)$. On the other hand, the union of these cells covers $R\mathbb{S}^{1}$, and the diameter of each cell is at most $2\hat{R}$. Thus, $\operatorname{card}(\mathcal{T}_{bd}(R))\geq\frac{2\pi}{2\arcsin\frac{\hat{R}}{R}}$, implying that $\operatorname{card}(\mathcal{T}_{bd}(R))=\Omega(R)$. ∎
@@
For any $R>2\hat{R}$, set $v_{R}(\mathcal{T})=\frac{\sum_{C\in\mathcal{T}(R)}v(C)}{\operatorname{card}(\mathcal{T}(R))}$. Then, $v_{R}(\mathcal{T})\leq 6+\mathcal{O}\left(\frac{1}{R}\right)$, and $\overline{v}(\mathcal{T})\leq 6$.
@@
Consider the graph $G$ whose vertices are the vertices of the cells of $\mathcal{T}(R)$, and two vertices are connected by an edge if they are the endpoints of an edge of a cell in $\mathcal{T}(R)$. Then $G$ is a planar graph. Let $V(G)$, $E(G)$ and $F(G)$ denote the numbers of vertices, edges and faces of $G$. Furthermore, let $\mathcal{E}_{bd}(G)$ denote the family of the edges of $G$ in the boundary of $\bigcup\mathcal{T}(R)$. We note that every edge in $\mathcal{E}_{bd}(G)$ belongs to a cell of $\mathcal{T}$ intersecting $R\mathbb{S}^{1}$, and thus, it follows from Lemmas 2.1 and 2.2 that $E_{bd}(G):=\operatorname{card}(\mathcal{E}_{bd}(G))=\mathcal{O}(R)$. These lemmas, an area estimate and the observation that $\bigcup\mathcal{T}(R)$ contains $(R-2\hat{R})\mathbf{B}^{2}$ and is contained in $(R+2\hat{R})\mathbf{B}^{2}$ shows that $V(G)$, $E(G)$ and $F(G)$ are all of order of magnitude $\Theta(R^{2})$.
We observe that there might be vertices of $G$ whose degree is less than $3$. If $p$ is such a vertex, then its degree is $2$, it lies in $\operatorname{bd}\left(\bigcup\mathcal{T}(R)\right)$, and it belongs to two edges in $\mathcal{E}_{bd}(G)$. Thus, the number of such edges is at most $E_{bd}(G)=\mathcal{O}(R)$. We merge these edges into one edge (i.e. we apply SP-reduction) and continue this process until we obtain another planar graph $G^{\prime}$ in which the degree of every vertex is at least $3$. If $V(G^{\prime}),E(G^{\prime})$ and $F(G^{\prime})$ denote the numbers of vertices, edges and faces of $G^{\prime}$, then by the construction of $G^{\prime}$ we have $V(G^{\prime})\leq V(G)$ and $E(G^{\prime})\leq E(G)$ with $V(G)-V(G^{\prime})=E(G)-E(G^{\prime})=\mathcal{O}(R)$, and $F(G^{\prime})=F(G)$.
Applying Euler’s formula for $G^{\prime}$ and using the fact that the degree of every vertex of $G^{\prime}$ is at least $3$ yields that $E(G^{\prime})\leq 3F(G^{\prime})-6$, implying that $E(G)\leq 3F(G)+\mathcal{O}(R)$. On the other hand, $v_{R}(\mathcal{T})\leq\frac{2E(G)}{F(G)}$, and as $E(G)$ and $F(G)$ are of $\Theta(R^{2})$, it follows that $v_{R}(\mathcal{T})\leq 6+\mathcal{O}\left(\frac{1}{R}\right)$, which readily yields the assertion. ∎
@@
Figure 1. The isoperimetrix $M_{iso}$ of a norm with unit disk $M$. The dotted circle is the Euclidean unit disk $\mathbf{B}^{2}$ centered at $o$. The left-hand side panel shows $M$ and its polar $M^{\circ}$, the isoperimetrix in the right-hand side panel is a rotated copy of $M^{\circ}$ by $\frac{\pi}{2}$
@@
A variant of this inequality, in Euclidean plane, states that among convex polygons with given perimeter and fixed directions of their sides, the ones circumscribed about a circular disk have maximal area. This statement is called L’huilier’s inequality, which is generalized for arbitrary normed planes by Chakerian [6] in the following way.
@@
Now we are ready to prove Theorem 1.9. For simplicity, we assume that $M_{iso}$ satisfies the $\alpha$-Dowker property for some $\alpha\in(0,\infty)$; for the case $\alpha=0$ a straightforward modification of our argument can be applied. From now on we also assume that every cell of $\mathcal{T}$ has unit area. We need to consider the following quantity.
$\underline{P}_{2\alpha}(\mathcal{T})=\liminf_{R\to\infty}\frac{\sum_{C\in\mathcal{T}(R)}\left(\operatorname{perim}_{M}(C)\right)^{2\alpha}}{\operatorname{card}(\mathcal{T}(R))}.$
@@
$\underline{P}_{2\alpha}(\mathcal{T})\geq 4^{\alpha}\liminf_{R\to\infty}\frac{\sum_{C\in\mathcal{T}(R)}\left(\operatorname{area}(C^{*})\right)^{\alpha}}{\operatorname{card}\mathcal{T}(R)}\geq 4^{\alpha}\liminf_{R\to\infty}\frac{\sum_{C\in\mathcal{T}(R)}\left(A_{M_{iso}}(v(C))\right)^{\alpha}}{\operatorname{card}\mathcal{T}(R)}.$
where $A_{M_{iso}}(v(C))$ denotes the area of a minimum area polygon with $v(C)$ sides, circumscribed about $M_{iso}$ (see Theorem 1.5). By our assumptions, the sequence $\{\left(A_{M_{iso}}(n)\right)^{\alpha}\}$ is convex, and by Lemma 2.4, $\overline{v}(\mathcal{T})\leq 6$. Thus,
$\underline{P}_{2\alpha}(\mathcal{T})\geq 4^{\alpha}\left(A_{M_{iso}}(6)\right)^{\alpha}.$
Now the only thing we need to show is that for some hexagonal tiling $\mathcal{T}_{hex}$ of $\mathcal{M}$, we have $P_{2\alpha}(\mathcal{T}_{hex})=4^{\alpha}\left(A_{M_{iso}}(6)\right)^{\alpha}$. To do it we recall another theorem of Dowker from [8].
@@
Let $K$ be an $o$-symmetric convex disk. Then, for every $m\geq 2$, there is an $o$-symmetric convex $(2m)$-gon $P$ circumscribed about $K$ with $\operatorname{area}(P)=A_{K}(2m)$.
By Theorem 2.7, there is an $o$-symmetric hexagon $H$ circumscribed about $M_{iso}$ of area $A_{M_{iso}}(6)$. Let $\mathcal{T}_{hex}$ be a tiling whose cells are translates of a homothetic copy of $H$ of unit area. Then the desired equality follows from the equality part of Theorem 2.6.
Remark 2.8.
If the isoperimetrix $M_{iso}$ of the norm has at most six sides (which is equivalent to the property that $M$ has at most six sides), then $M$ satisfies the log-honeycomb property.